The reaction A(aq) → B(aq) is a first order reaction. If it takes 300.0 seconds for the concentration of A to decrease in concentration from 0.50 M to 0.010 M, what is the rate constant for this reaction?

57 s-1
1.5 x 10-2 s-1
1.3 x 10-2 s-1
1.6 x 10-3 s-1
1.8 x 10-2
Answer #1

Solution:

For a first order reaction, the expession for rate constant (k) is expressed as:

k = (2.303/t) log (a/a-x)

a = initial concentration = 0.50 M

(a-x) =final concentration = 0.010 M

t = time = 300 sec, then,

k = (2.303/ 300 sec) log (0.50 M / 0.010 M)

k = (2.303/ 300 sec) log 50

k = (2.303 / 300 sec) x 1.6990

k = 1.3 x 10^-2 sec-1

Hence, option is C.